x+1/2 - x+2/3 = x+3/4
solve for x

Answer by checkley75(3666)

x+1/2-x+2/3=x+3/4
x-x-x=3/4-1/2-2/3
-x=(9-6-8)/12
-x=5/12
x=-5/12 answer.